What is the probability of finding a n = 2, ℓ = 1 electron between a0 and 2a0?
1
Expert's answer
2020-06-17T09:27:39-0400

The probability that such electron is in the region a0<r<2a0 is


"P=\int^{2a_0}_{a_0}\frac{r^4}{24a_0^5}e^{-r/a_0}\text{d}r=0.04899."